OpenStudy (rainbowdashrocks):

Jacob’s recipe calls for 4 7/8 cups of raisins. He has 6/11 of the amount that he needs. About how many cups of raisins does Jacob have? Estimate by first rounding each number to the nearest 1/2 . A.4 cups B.5 1/2 C.2 1/2 D.3 1/2 PLZ HELP WILL FAN AND MEDAL

OpenStudy (rainbowdashrocks):

help plz

OpenStudy (escamer):

my best guess would be D but im not 100% sure on that, ill tag someone that will know the answer @malcolmmcswain @the_ocean_girl

Theoceangirl (the_ocean_girl):

First we need to make 4 7/8 an improper fraction

OpenStudy (rainbowdashrocks):

ok

OpenStudy (rainbowdashrocks):

39/8

OpenStudy (malcolmmcswain):

Ocean beat me D:

OpenStudy (rainbowdashrocks):

lol

Theoceangirl (the_ocean_girl):

do you know how to do that?

OpenStudy (rainbowdashrocks):

39/8 is my improper fraction

Theoceangirl (the_ocean_girl):

Good! Now when you divide by fractions you have to flip it and the sign. So it would be 39/8 * 11/6

OpenStudy (rainbowdashrocks):

cross multiply?

OpenStudy (rainbowdashrocks):

429/48?

OpenStudy (rainbowdashrocks):

hello?

OpenStudy (escamer):

@YoungStudier ocean_girl had to go do something but this person can help you

OpenStudy (rainbowdashrocks):

ok

OpenStudy (rainbowdashrocks):

i think c

OpenStudy (rainbowdashrocks):

anyone here i feel i am texting a wall

OpenStudy (youngstudier):

@malcolmmcswain

OpenStudy (escamer):

mal is doing his college homework young ._.

OpenStudy (thadds2003):



OpenStudy (rainbowdashrocks):

lol

OpenStudy (thadds2003):



OpenStudy (rainbowdashrocks):

i need help though its my last q

OpenStudy (rainbowdashrocks):

@ItsKoreyyy

OpenStudy (rainbowdashrocks):

i need help

OpenStudy (anonymous):

I can try! :P Could you open this up in a new question tab?

OpenStudy (rainbowdashrocks):

dont know how

OpenStudy (rainbowdashrocks):

i think its c but i am not sure

OpenStudy (michele_laino):

hint: your exercise can be modeled by this computation: \[\frac{6}{{11}} \times \left( {4\frac{7}{8}} \right) = \frac{6}{{11}} \times \left( {4 + \frac{7}{8}} \right) = ...?\]

OpenStudy (rainbowdashrocks):

i got 429/48

OpenStudy (rainbowdashrocks):

is that right